A firm is producing 100 units of its product. At this level of output the firm’s average variable
cost AVC = $80, and the average total cost ATC = $120. The firm is a price taker and the price
for its product is $100. Assuming that the firm is maximizing profits and that labour is the only
variable input, answer the following questions: a. What are the fixed costs of the firm? Give a number.
b. What are the profits or losses?
c. What are the rents, if any?
d. Is the firm producing at the minimum ATC? Explain.
e. Does the firm experience diminishing marginal products of labor at this level of output? explain
1
Expert's answer
2020-03-03T09:33:29-0500

a. "FC=AFC *Q"

FC= 40*100= 4000$

b. TC>VC = loss

Loss =12000-8000 =4000$

c. Rent=marginal cost -opportunity cost.

Here no opportunity cost involved hence no rent

d. At minimum ATC;"ATC =mc"

Mc =12000 /100=120$

ATC =120$

e.There are diminishing marginal products of labour at this level of output because the firm makes losses